OpenStudy (anonymous):

find the exact value of sin2theta given , Costheta=1/6,0degrees,

OpenStudy (campbell_st):

well use the information  find x using pythagoras' theorem then you can find sin(theta) then use the expansion \[\sin(2 \theta) = 2 \sin(\theta)\cos(\theta)\]